OpenStudy (anonymous):

how to find sec(3pi\4)

OpenStudy (anonymous):

\[\sec(\frac{3\pi}{4})=\frac{1}{\cos(\frac{3\pi}{4})}\]

OpenStudy (anonymous):

sec is 1/cos so u go to 3pi/4 and look at the x coordinate which is -sqrt2/2 nd since its under 1 u basically flip it givving u -2/sqrt2 then u multiply top and bottom by sqrt2 to get rid of radical from bottom so you would get -2sqrt2/2 finallt u can cancel the 2s and your final answer would be -sqrt2